OpenStudy (anonymous):

WILL FAN AND GIVE MEDAL TO ANYONE WHO CAN ANSWER THIS CORRECTLY! Rita and Jackie are raking leaves to earn some money. Rita can rake 1 lawn in 3 hours. Jackie can rake 1 lawn in 4 hours. How long would it take both girls to work together to rake 1 lawn? Write and solve an equation for this situation. Explain how to set up the equation, using w = rt.

OpenStudy (anonymous):

R = 1/3 lawn/hour J = 1/4 lawn/hour R+J = 1/3 + 1/4 lawn/hour = 7/12 lawn per hour Since together they can rake 7/12 lawn per hour, it would take them 12/7 hours = 1 5/7 hours ≈ 1 hour 43 minutes to rake one lawn